# Binary Autoencoder for Mechanistic Interpretability of Large Language Models

### v1

@@ Abstract
Existing works are dedicated to untangling atomized numerical components (features) from the hidden states of Large Language Models (LLMs) for interpreting their mechanism. However, they typically rely on autoencoders constrained by some implicit training-time regularization on single training instances (i.e., $L_{1}$ normalization, top-k function, etc.), without an explicit guarantee of global sparsity among instances, causing a large amount of dense (simultaneously inactive) features, harming the feature sparsity and atomization. In this paper, we propose a novel autoencoder variant that enforces minimal entropy on minibatches of hidden activations, thereby promoting feature independence and sparsity across instances. For efficient entropy calculation, we discretize the hidden activations to 1-bit via a step function and apply gradient estimation to enable backpropagation, so that we term it as Binary Autoencoder (BAE) and empirically demonstrate two major applications: (1) Feature set entropy calculation. Entropy can be reliably estimated on binary hidden activations, which we empirically evaluate and leverage to characterize the inference dynamics of LLMs and In-context Learning. (2) Feature untangling. Similar to typical methods, BAE can extract atomized features from LLM’s hidden states. To robustly evaluate such feature extraction capability, we refine traditional feature-interpretation methods to avoid unreliable handling of numerical tokens, and show that BAE avoids dense features while producing the largest number of interpretable ones among baselines, which confirms the effectiveness of BAE serving as a feature extractor [^1] The source code and the BAE checkpoints will be released upon acceptance of this paper..
@@ 1 Introduction
Current practice for untangling atomized numerical components (features) from Large Language Models (LLMs), such as Sparse Autoencoder (SAE) (Shu et al., 2025a), with training-time regularization (e.g., $L_{1}$ normalization on hidden activations) to implicitly atomize features sample-wisely. However, such methodologies do not ensure global sparsity, often leading to frequently activated (dense) features alongside inactive (dead) features (Stolfo et al., 2025; Rajamanoharan et al., 2024b; Sun et al., 2025), contradicting the sparsity assumption (Elhage et al., 2022) in mechanistic interpretability, as broad activations across samples hinder consistent and meaningful interpretation, and reduce the parameter efficiency by dead features.
@@
Therefore, in this paper, to address this issue, we propose Binary Autoencoder (BAE), utilizing information-theoretic constraints among minibatches of training instances to address the aforementioned issue. As shown in Fig. 1, we design training objectives that minimize the entropy of hidden activation from minibatches to reduce feature covariation while enforcing global sparsity to suppress frequently activated features. However, typical hidden activations are real vectors, whose entropies are extremely difficult to calculate (Greenewald et al., 2023) as the computation requires high-dimensional numerical integrations, which involve an exponential explosion of computational complexity. To this end, we round the activations to binary, and calculate entropy on such binary vectors to significantly reduce the computational complexity of the entropy objective and utilize gradient estimation (Hubara et al., 2016) to enable the backpropagation on such a rounding operation.
@@
We empirically demonstrated the benefits of BAE in application as: (1) Efficient estimation of entropy from hidden state sets. We can estimate the entropy for reconstructing the input sets as the hidden activation entropy with significantly reduced cost, which can serve as an important metric for understang the inner operation of LLMs. By experiment on synthetic datasets with different ground-truth entropy values, we confirm the accuracy of such entropy estimation. Moreover, we investigate the entropy dynamics inherent in the feed-forward process of standard language modeling in LLMs, revealing an information bandwidth at each layer and implicit context windows. Also, we further interpret In-context Learning (ICL) (Dong et al., 2022) as a form of information reduction.
And (2) Sparse feature extraction. As also applied to normal SAE, the line vectors of the second autoencoder layer (sometimes termed as the dictionary) can serve as the atomized features extracted from the inputs. In typical SAE, features are activated densely (Stolfo et al., 2025; Rajamanoharan et al., 2024b; Sun et al., 2025) due to the inconsistent scales across hidden activation channels, rooted in the sample-wise regularization, which overlooks the global sparsity. In contrast, BAE significantly suppresses dense features as well as dead features by the entropy-based training objective on mini-batches and consistent activation scaling based on the information gain of each channel, extracting the largest number of active and interpretable features. Meanwhile, we refine the traditional feature interpretation method (Bills et al., 2023; Huben et al., 2024), thereby avoiding unreliable handling of numerical tokens by LLMs (Press et al., 2023), for more robust automatic feature interpretation.
@@
Main findings and contributions: (1, §3) We propose Binary Autoencoder (BAE), a novel variant of autoencoder with binarized hidden activation and training-time entropy constraint. (2, §4) BAE enables accurate entropy calculation on hidden state sets, providing an important tool for analyzing LLM behavior from an information perspective. With this, we observe LLM inference as an information increment: hidden states at each layer act as channels with increasing bandwidth w.r.t. the layer depth, where each token adds information until the limit is reached, while ICL appears as information reduction. (3, §5) We confirm BAE as an effective atomized feature extractor with no dense features and the largest amount of interpretable features extracted. (4, §5) We revise the typical automatic feature interpretation (Bills et al., 2023; Huben et al., 2024) methods working on the line vector of the second BAE/SAE layer, which forces LLMs to process numerical tokens unreliably (Press et al., 2023), for interpreting the features extracted by BAE more robustly.
@@ 2 Related Works
(Shu et al., 2025b). Modern mechanistic interpretability (Sharkey et al., 2025; Bereska & Gavves, 2024) views LLM hidden states as superpositions of atomic features, motivating efforts to disentangle them for better semantical understanding of the LLMs’ operation. Originally, Sparse Autoencoder (SAE), an autoencoder with training-time $L_{1}$ normalization on the hidden activations for the sample-wise sparsity, is utilized for such disentangling (Bricken et al., 2023; Huben et al., 2024; Templeton et al., 2024; Gao et al., 2025). Additionally, variants of SAE are proposed to enhance the performance by utilizing modified activation functions (Gao et al., 2025; Bussmann et al., 2024; Rajamanoharan et al., 2024a; b) or merging multiple-layer hidden states from LLMs (Shi et al., 2025; Lindsey et al., 2024). However, these methods do not warrant a global sparse decomposition of hidden states, that is, with some guarantee to minimize the activation frequency as well as covariation among channels of hidden activations to reconstruct the original hidden states with sparse atomized features. Such a drawback causes the activations to be dense (Kissane et al., 2024; Rajamanoharan et al., 2024b; Sun et al., 2025; Stolfo et al., 2025), i.e., some features are activated over diverse inputs lacking a common semantic theme, and thus cannot be uniquely and clearly attributed (or interpreted) semantically. Therefore, these dense features can be regarded as a merger of multiple features, i.e., insufficient atomization, which is the main motivation of this paper: explicitly promote the global feature sparsity and atomization by improved training methods, to capture a larger amount of interpretable features, and to improve the accuracy of feature interpretation.
@@
To utilize the information-theoretic constraints mentioned before, we round the hidden activations to binary. The framework of such binary neural networks, including rounding the real activation into 1-bit, and estimating the gradient for such rounding operation, etc., is originally proposed by Hubara et al. (2016). Subsequently, numerous variants are proposed primarily based on different rounding operations and gradient estimations (Rastegari et al., 2016; Zhou et al., 2016; Choi et al., 2018; Vargas et al., 2024). A detailed survey can be found in Qin et al. (2020). In this paper, we utilize these methodologies for the autoencoder with binarized hidden activations to reduce the calculation cost on the information-theoretic constraints.
Figure 1: Feed-forward computation and training objective of BAE. Hidden states $h_{0}$ from LLM layers are mapped by $W_{\text{in}}$, binarized into $h_{1}$ via a step function, and projected back by $W_{\text{out}}$ as $\hat{h_{0}}$. The $\hat{h_{0}}$ is fed to the self-regression loss, while $h_{1}$ is fed to the information bottleneck loss.
@@ 3 Binary Autoencoder
To address SAE’s drawbacks, we propose Binary Autoencoder (BAE) shown in Fig. 1 and follows:
@@
Given input hidden state $h_{0}\in\mathbb{R}^{d}$ from an LLM, a binary autoencoder $\mathcal{F}$ conducts such calculation for the output $\mathcal{F}(h_{0})$ (also noted as $\hat{h_{0}}=\mathcal{F}(h_{0})$):
@@
where $W_{\text{in}}\in\mathbb{R}^{d\times d^{\prime}}$, preforming a linear decomposition on input $h_{0}$ to $d^{\prime}$ dimensions, $b\in\mathbb{R}^{d}$ is a bias term to reconstruct the anisotropism (Gao et al., 2019; Ethayarajh, 2019; Biś et al., 2021; Godey et al., 2024) of the hidden states which contains minor information thus should be ignored as background value, $\Gamma$ is the quantization function, projecting $\mathbb{R}^{d^{\prime}}$ into $\{0,1\}^{d^{\prime}}$ element-wisely. Here, we use the step function to binarize the $h_{0}W_{\text{in}}$ into hidden activation $h_{1}$:
$\Gamma\left(\left[x_{1},x_{2},\dots,x_{d^{\prime}}\right]\right)$ $=\left[\gamma\left(x_{1}\right),\gamma\left(x_{2}\right),\dots,\gamma\left(x_{d^{\prime}}\right)\right],$
@@
The $\Gamma$ also provides essential non-linearity for the numerical expressivity of Eq. 1 (otherwise, the $W_{\text{in}}W_{\text{out}}$ will degrade into one matrix).
Self-regression Training Objective.
@@
Information Bottleneck Constraint.
To constrain the hidden activation ($h_{1}=\Gamma\left(h_{0}W_{\text{in}}\right)$) to a global sparse representation for $h_{0}$, we minimize the margin entropy of $h_{1}$ (notice that, since $h_{1}\in\{0,1\}^{d^{\prime}}$, such calculation are differentiable and highly efficient, without any numerical integration on the real space). Also, to maximize the effectiveness of the constraint on margin entropy, we also penalize the covariance (except the diagonal elements) of $h_{1}$ to force the margin entropy to approach the joint distribution entropy. That is, on the minibatch $H_{0}$, we define the entropy-based loss term:
@@
The $\alpha_{e}$ and $\alpha_{c}$ are hyperparameters. The total loss can be written as:
@@
This training objective resembles and similates the information bottleneck methods (Tishby et al., 2000; Kawaguchi et al., 2023; Tishby & Zaslavsky, 2015), where $\mathcal{L}_{e}$ minimizes the mutual information between input $h_{0}$ and latent $h_{1}$, while $\mathcal{L}_{r}$ maximizes that between output and $h_{1}$.
@@
Since the differential of binarization function $\Gamma$ is 0 almost everywhere, to enable the backpropagation from $\mathcal{L}$ to $W_{\text{in}}$, following the previous works (Hubara et al., 2016; Vargas et al., 2024), we estimate the gradient of $\Gamma$ by smoothing function $x\mapsto(1+e^{-x})^{-1}$ (“sigmoid”) elementwisely. Therefore, the differential of $\Gamma$ is estimated as:
$\frac{\partial\Gamma(x)}{\partial x}\coloneqq x\odot(\mathbf{1}-x),$ (6)
where $\mathbf{1}$ is all-ones vector, $\odot$ is the Hadamard product of two vectors.
@@
We defaultly set $d^{\prime}=4d$ (i.e., the expanding rate =4=4), the entropy loss weight [^2] Notice that, considering the magnitude around the late training stage (as shown in Fig. 2 and 3) of the self-regression loss ($\mathcal{L}_{r}$, around $10^{-3}\sim 10^{-2}$) and the margin entropy ($\mathrm{H}[\cdot]$, around $10^{0}\sim 10^{3}$), athough the weights $\alpha_{e}$ are set quite small, the magnitude of the $\alpha_{e}\mathrm{H}[\cdot]$ are balanced as a regularization term against $\mathcal{L}_{r}$. $\alpha_{e}=10^{-7}$, the covarience loss weight $\alpha_{c}=10^{-7}$. We use Adam (Kingma & Ba, 2014) optimizer with learning rate $5\times 10^{-4}$, momentum factor $\alpha_{1}=0.9$, $\alpha_{2}=0.999$, and minibatch size $n_{b}=512$ for $2000$ epochs, with $\alpha_{e}=0$ in the first $500$ epochs.
4 Entropy Estimation of Hidden States by BAE
@@
As a direct measurement of information amount, calculating entropy for hidden states in neural networks can promote a closer observation into the inner mechanism. However, directly computing the differential entropy of a high-dimensional vector set requires probability density estimation (DE) and integration, which is neither accurate (DE suffers from the curse of dimensionality and floating-point errors) nor efficient (numerical integration costs $\mathrm{O}(C^{d})$ for $d$ dimensions with $C$ integration cells) (Greenewald et al., 2023). In contrast, BAE can improve such entropy calculation by best-effortly decomposing the original real-value vectors into binary vectors ($h_{1}$) with pair-wisely independent elements, enabling efficient entropy estimation by margin entropy of the mean hidden activation $\bar{h_{1}}$: given a vector set $H_{0}=\left\{h_{0}^{(i)}\right\}_{i=1}^{n}$, we encode them into $h_{1}$s by a trained BAE as $H_{1}=\left\{h_{1}^{(i)}=\Gamma\left(h_{0}^{(i)}W_{\text{in}}\right)\right\}_{i=1}^{n}$. Since the $h_{1}$s are binarized and best-effortly decorrelated pairwise by the covariance loss, we can calculate the entropy required for reconstructing $H_{0}$ as $\mathrm{H}[\bar{h_{1}}]$, where $\bar{h_{1}}$ are the averages among the line vectors of $H_{1}$.
In this section, we evaluate the aforementioned entropy estimation on a synthesis dataset to show its accuracy, and utilize such entropy calculation to track the feed-forward process of LLM.
4.1 Evaluating Entropy Calculation from Binary Autoencoder

### v2

@@ Abstract
Existing works are dedicated to untangling atomized numerical components (features) from the hidden states of Large Language Models (LLMs). However, they typically rely on autoencoders constrained by some training-time regularization on single training instances, without an explicit guarantee of global sparsity among instances, causing a large amount of dense (simultaneously inactive) features, harming the feature sparsity and atomization. In this paper, we propose a novel autoencoder variant that enforces minimal entropy on minibatches of hidden activations, thereby promoting feature independence and sparsity across instances. For efficient entropy calculation, we discretize the hidden activations to 1-bit via a step function and apply gradient estimation to enable backpropagation, so that we term it as Binary Autoencoder (BAE) and empirically demonstrate two major applications: (1) Feature set entropy calculation. Entropy can be reliably estimated on binary hidden activations, which can be leveraged to characterize the inference dynamics of LLMs. (2) Feature untangling. Compared to typical methods, due to improved training strategy, BAE avoids dense features while producing the largest number of interpretable ones among baselines.
@@ 1 Introduction
Current practice for untangling atomized numerical components (features) from Large Language Models (LLMs), such as Sparse Autoencoder (SAE) (shu2025survey), applies training-time regularization (e.g., $L_{1}$ normalization on hidden activations) to implicitly atomize features sample-wisely. However, such methodologies do not ensure global sparsity, often leading to frequently activated (dense) features alongside inactive (dead) features (stolfo2025antipodal; rajamanoharan2024jumping; sun2025dense), contradicting the sparsity assumption (elhage2022toymodelsof) of LLM hidden states, as broad activations across samples hinder consistent and meaningful interpretation, and reduce the parameter efficiency by dead features.
Figure 1: Feed-forward computation and training objective of BAE. Hidden states $h_{0}$ from LLM layers are mapped by $W_{\text{in}}$, binarized into $h_{B}$ via a step function, and projected back by $W_{\text{out}}$ as $\hat{h_{0}}$. The $\hat{h_{0}}$ is fed to the self-regression loss, while $h_{B}$ is fed to the information bottleneck loss. More-bit decode: to reduce the information loss of the BAE, as mentioned in §6, we aggregate real-valued hidden activation elements from multiple binary bits, and perform decoding using the reconstructed real-valued vector.
@@
Therefore, in this paper, to address this issue, we propose Binary Autoencoder (BAE), utilizing information-theoretic constraints among minibatches of training instances to address the aforementioned issue. As shown in Fig. 1, we design training objectives that minimize the entropy of hidden activation from minibatches to reduce feature covariation while enforcing global sparsity to suppress frequently activated features. However, typical hidden activations are real vectors, whose entropies are extremely difficult to calculate (greenewald2023high) as the computation requires high-dimensional numerical integrations, which involve an exponential explosion of computational complexity. To this end, we round the activations to binary, then calculate entropy on such binary vectors to significantly reduce the computational complexity of the entropy objective, and utilize gradient estimation (hubara2016binarized) to enable the backpropagation on such a rounding operation.
@@
We empirically demonstrated the benefits of BAE in application as: (1) Efficient estimation of entropy from hidden state sets. We can estimate the entropy for reconstructing the input sets as the hidden activation entropy with significantly reduced cost. By experiment on synthetic datasets with different ground-truth entropy values, we confirm the accuracy of such entropy estimation. Moreover, we utilize entropy to track the feed-forward process of standard language modeling in LLMs, revealing an information bandwidth at each layer and implicit context windows. Also, we further interpret In-context Learning (ICL) (dong2022survey) as a form of information reduction. And (2) Sparse feature extraction. As also applied to normal SAE, the row vectors of the second autoencoder layer (sometimes termed as the dictionary) can serve as the atomized features extracted from the inputs. Compared to typical SAE, BAE significantly suppresses dense features as well as dead features by the entropy-based training objective on mini-batches and consistent activation scaling based on the information gain of each channel, extracting the largest number of active and interpretable features. (3) More-bit decoding. Also, to reduce the information loss induced by the rounding operation, we adopt a more-bit decoding as a handle for the trade-off between feature atomization and information loss.
@@
Main findings and contributions: (1, §3) We propose Binary Autoencoder (BAE), a novel variant of autoencoder with binarized hidden activation and training-time entropy constraint. (2, §4) BAE enables accurate entropy calculation on hidden state sets, and we utilize it for analyzing LLM behavior from an information perspective. (3, §5) We confirm BAE as an effective atomized feature extractor with no dense features and the largest amount of features extracted.
@@ 2 Related Works
(shu2025survey). Modern mechanistic interpretability (sharkey2025open; bereska2024mechanistic) views LLM hidden states as superpositions of atomic features, motivating efforts to disentangle them for better semantical understanding of the LLMs’ operation, where Sparse Autoencoder (SAE), autoencoder with training-time $L_{1}$ normalization on the hidden activations for the sample-wise sparsity, and its variants (gao2025scaling; bussmann2024batchtopk; rajamanoharan2024improving; rajamanoharan2024jumping; shi2025route; crosscoder) is applied (bricken2023monosemanticity; huben2024sparse; templeton2024scaling; gao2025scaling). However, these methods do not warrant a global sparse disentanglement, that is, with some guarantee to minimize the activation frequency as well as covariation among channels of hidden activations. Such a drawback causes the activations to be dense (kissane2024interpreting; rajamanoharan2024jumping; sun2025dense; stolfo2025antipodal), i.e., some features are activated across diverse inputs without a shared semantics, and thus cannot be uniquely interpreted. Moreover, these dense features can be viewed as mergers of multiple features (i.e., insufficient atomization), which is the main motivation of this paper: explicitly promote the global feature sparsity and atomization by improved training methods, to capture a larger amount of interpretable features, and to improve the accuracy of feature interpretation.
@@
To utilize the information-theoretic constraints mentioned before, we round the hidden activations to binary. The framework of such binary neural networks, including rounding the real activation into 1-bit, and estimating the gradient for such rounding operation, etc., is originally proposed by (hubara2016binarized). Subsequently, numerous variants are proposed primarily based on different rounding operations and gradient estimations (rastegari2016xnor; zhou2016dorefa; choi2018pact; vargas2024biper). A detailed survey can be found in (qin2020binary). In this paper, we utilize these methodologies to reduce the calculation cost on the information-theoretic constraints.
@@ 3 Binary Autoencoder
To address SAE’s drawbacks, we propose Binary Autoencoder (BAE) shown in Fig. 1 and follows: Given input hidden state $h_{0}\in\mathbb{R}^{d}$ from an LLM, a binary autoencoder $\mathcal{F}$ conducts such calculation for the output $\mathcal{F}(h_{0})$ (also noted as $\hat{h_{0}}=\mathcal{F}(h_{0})$) as:
@@
where $W_{\text{in}}\in\mathbb{R}^{d\times d^{\prime}}$, preforming a linear decomposition on input $h_{0}$ to $d^{\prime}$ dimensions, $b\in\mathbb{R}^{d}$ is a bias term reconstructing the hidden states anisotropism (gao2018representation; ethayarajh-2019-contextual; bis-etal-2021-much; godey-etal-2024-anisotropy) which contains minor information thus should be ignored as background value, $\Gamma$ is the quantization function, projecting $\mathbb{R}^{d^{\prime}}$ into $\{0,1\}^{d^{\prime}}$ element-wisely. Here, we use the step function to binarize the $h_{0}W_{\text{in}}$ into hidden activation $h_{B}$:
$h_{B}=\Gamma\left(\left[x_{1},x_{2},\dots,x_{d^{\prime}}\right]\right)$ $=\left[\gamma\left(x_{1}\right),\gamma\left(x_{2}\right),\dots,\gamma\left(x_{d^{\prime}}\right)\right],$
@@
The $\Gamma$ also provides essential non-linearity for the numerical expressivity of the autoencoder (otherwise, the $W_{\text{in}}W_{\text{out}}$ will degrade into one matrix).
Self-regression Training Loss.
@@
Information Bottleneck (Entropy) Loss.
To constrain the hidden activation ($h_{B}=\Gamma\left(h_{0}W_{\text{in}}\right)$) to a global sparse representation for $h_{0}$, we minimize the margin entropy of $h_{B}$ (since $h_{B}\in\{0,1\}^{d^{\prime}}$, such calculation are differentiable and efficient, without any numerical integration on the real space). Also, to maximize the effectiveness of the constraint on margin entropy, we also penalize the covariance (except the diagonal elements) of $h_{B}$ to force the margin entropy to approach the joint distribution entropy. That is, on the minibatch $H_{0}$, we define the entropy-based loss term:
@@
The $\alpha_{e}$ and $\alpha_{c}$ are hyperparameters. The total loss is:
@@
This training objective resembles and simulates the information bottleneck methods (tishby2000information; kawaguchi2023does; tishby2015deep), where $\mathcal{L}_{e}$ minimizes the mutual information between input $h_{0}$ and latent $h_{B}$, while $\mathcal{L}_{r}$ maximizes that between output and $h_{B}$.
@@
Since the differential of binarization function $\Gamma$ is 0 almost everywhere, to enable the backpropagation from $\mathcal{L}_{e}$ to $W_{\text{in}}$, following the previous works (hubara2016binarized; vargas2024biper), we estimate the gradient of $\Gamma$ by smoothing function $x\mapsto(1+e^{-x})^{-1}$ (“Sigmoid”) elementwisely:
$\frac{\partial\Gamma(x)}{\partial x}\coloneqq\mathrm{Sigmoid}(x)\odot(\mathbf{1}-\mathrm{Sigmoid}(x)),$ (6)
where $\mathbf{1}$ is all-ones vector, $\odot$ is the Hadamard product.
@@
We defaultly set $d^{\prime}=4d$ (i.e., the expanding rate =4=4), the entropy loss weight [^1] Notice that, considering the magnitude around the late training stage (as shown in Fig. 2 and 12) of the self-regression loss ($\mathcal{L}_{r}$, around $10^{-3}\sim 10^{-2}$) and the margin entropy ($\mathrm{H}[\cdot]$, around $10^{0}\sim 10^{3}$), athough the weights $\alpha_{e}$ are set quite small, the magnitude of the $\alpha_{e}\mathrm{H}[\cdot]$ are balanced as a regularization term against $\mathcal{L}_{r}$. $\alpha_{e}=10^{-7}$, the covarience loss weight $\alpha_{c}=10^{-7}$. We use Adam (kingma2014adam) optimizer with learning rate $5\times 10^{-4}$, momentum factor $\alpha_{1}=0.9$, $\alpha_{2}=0.999$, and minibatch size $n_{b}=512$ for $2000$ epochs, with $\alpha_{e}=0$ in the first $500$ epochs.
4 Entropy Estimation of Hidden States
@@
As a direct measurement of information amount, calculating entropy for hidden states in neural networks can promote a closer observation into the inner mechanism. However, directly computing the differential entropy of a high-dimensional vector set requires probability density estimation (DE) and integration, which is neither accurate (DE suffers from the curse of dimensionality and floating-point errors) nor efficient (numerical integration costs $\mathrm{O}(C^{d})$ for $d$ dimensions with $C$ integration cells) (greenewald2023high). In contrast, BAE can improve such entropy calculation by best-effortly decomposing the original real-value vectors into binary vectors ($h_{B}$) with pair-wisely independent elements, enabling efficient entropy estimation by margin entropy of the mean hidden activation $\bar{h_{B}}$: given a vector set $H_{0}=\left\{h_{0}^{(i)}\right\}_{i=1}^{n}$, we encode them into $h_{B}$s by a trained BAE as $H_{B}=\left\{h_{B}^{(i)}=\Gamma\left(h_{0}^{(i)}W_{\text{in}}\right)\right\}_{i=1}^{n}$. Since the $h_{B}$s are binarized and best-effortly decorrelated pairwise by the covariance loss, we can calculate the entropy required for reconstructing $H_{0}$ as $\mathrm{H}[\bar{h_{B}}]$, where $\bar{h_{B}}$ are the averages among the row vectors of $H_{B}$.
In this section, we evaluate the aforementioned entropy estimation with best effort on a synthesis dataset to show its accuracy, and utilize such entropy calculation to track the feed-forward process of LLM.
4.1 Evaluating Entropy Calculation from BAE